OpenStudy (trevor12):

Which of the following numbers is rational? -9 0 all of the above

OpenStudy (18turns):

think what does ratinal mean

OpenStudy (solomonzelman):

Natural Numbers: \(\tiny \\[0.5em]\) \(\color{#000000 }{ \displaystyle 1,~~2,~~3,~~4,~~5,~~6,~~~~{\bf ...~}{\rm so~on~forever} }\) Whole Numbers: \(\tiny \\[0.5em]\) \(\color{#000000 }{ \displaystyle 0,~~1,~~2,~~3,~~4,~~5,~~6,~~~~{\bf ...~}{\rm so~on~forever} }\) (You can see that the whole numbers conatin "0", but the rest of them are the same.) Integers \(\tiny \\[0.5em]\) \(\color{#000000 }{ \displaystyle {\rm \longleftarrow}~{\bf .....~}-3,~~-2,~~-1,~~~~0,~~~~1,~~~~2,~~~~3,~{\bf .....~}{\rm \longrightarrow} }\) (Extends both ways. So -1023 or 600 (for instance are also integers). \(\color{#0fd0f0 }{ \displaystyle \bf (See~~that~~each~~set~~adds~~on~~more~~to~~previous?) }\)

OpenStudy (solomonzelman):

And, Rational Numbers: \(\tiny \\[0.5em]\) are numbers that can be recorded as a ratio of p/q, where p and q are both integers.

OpenStudy (solomonzelman):

For example, 9=9/1 =integer/integer 0.4/0.99=40/99 =integer/integer

OpenStudy (solomonzelman):

Also, -5.2 = -52/10 = integer/integer

OpenStudy (solomonzelman):

So those numbers are rationals